OVHCloud Web Hosting Status

Current status
  • Operational
  • Degraded performance
  • Partial Outage
  • Major Outage
  • Under maintenance
FS#4095 —private sql-chassis change
Scheduled Maintenance Report for Web Cloud
During the next two nights, we will change the chassis of current private sql and switch to a configuration which is:
. more efficient
. less sensitive to heat

This configuration is being tested for several months and the results are conclusive.

The migration will take place at night, and no data loss will occur. During this time, operations via manager/soap will be stopped.
For a few minutes, private sql will be in a \"read only\" mode, then it will be unavailable during the chassis change.


Date: 2010-05-04 09:00:37 UTC
server9 restrated, everything is functional

Date: 2010-05-04 09:00:15 UTC
Transfer completed, restarting of the server

Date: 2010-05-04 08:59:58 UTC
Stopping of server9

Date: 2010-05-04 08:59:31 UTC
We are going now to transfer server9 and server10

Date: 2010-05-04 08:59:06 UTC
server8 is ok

Date: 2010-05-04 08:58:55 UTC
Transfer of server8 completed, reboot of private sql is ongoing

Date: 2010-05-04 08:58:26 UTC
The private sqls start again on server6

Date: 2010-05-04 08:57:54 UTC
extinction of server6

Date: 2010-05-04 08:57:41 UTC
transfer of server5 completed, we reboot it

Date: 2010-05-04 08:57:16 UTC
We turn off server5

Date: 2010-05-04 08:56:42 UTC
Transfer completed, we reboot server 15

Date: 2010-05-04 08:56:26 UTC
Around server 15, we turn it off and we transfer data to the new chassis

Date: 2010-05-04 08:56:03 UTC
Data transfer completed for server4.
We reboot the private sqls

Date: 2010-05-04 08:55:38 UTC
We start this night with server4. Stopping of private sqls in a few minutes

Date: 2010-05-04 08:55:15 UTC
The rest will be on next night..

Date: 2010-05-04 08:53:59 UTC
Private sqls are restarted, everything is ok on server3

Date: 2010-05-04 08:53:38 UTC
data import completed, reboot is ongoing

Date: 2010-05-04 08:53:08 UTC
Around server3, a stopping is ongoing

Date: 2010-05-04 08:50:29 UTC
server2 is switched, reboot is ongoing

Date: 2010-05-04 08:50:11 UTC
The intervention is ok for server1, everything has been restarted and is functional.
We pass to server2, a stopping in a few minutes

Date: 2010-05-04 08:48:03 UTC
Private sqls of server1 restart, verification of good functioning...

Date: 2010-05-04 08:46:28 UTC
Chassis change completed, data import

Date: 2010-05-04 08:45:31 UTC
We start with server 1. A stopping for a few minutes.
Posted May 04, 2010 - 08:44 UTC